### v1

@@ 1 INTRODUCTION
Stray light from bright body or optical path dust reflections can also result in false positives. Given that the intensity of stray light often surpasses the background brightness level, conventional threshold-based star detection algorithms [5][7][8][9] are unable to discriminate false detections, undermining subsequent star identification processes. To counteract stray light, large baffles are frequently placed in front of the sensor optics [10], potentially encroaching on the limited payload space available in CubeSats. Even with a baffle, star trackers remain susceptible to blinding when directly exposed to bright bodies in their field of view [11].
@@
    Introducing an end-to-end training approach for a singular CNN, designed to concurrently handle star detection and centroiding. This method surpasses the centroiding accuracy and detection robustness achieved by several existing techniques.
@@
    Generating synthetic star images, augmented with real sensor noise and stray light, for training purposes. This approach significantly reduces the reliance on manual labeling efforts.
@@
    Conducting a comprehensive performance evaluation of various CNN models and identifying the most suitable CNN architectures for real-time star tracker image processing.
@@ 3.1 Training Data Generation
where the coordinates of the real star centroid are $u_{c}$, $v_{c}$. $\sigma_{\text{psf},u}$ and $\sigma_{\text{psf},v}$ are the PSF width in width and height direction in pixels. We choose $\sigma_{\text{psf},u}=\sigma_{\text{psf},v}=\sigma_{\text{psf}}$ in our simulation. In our experiments, we established that uniformly sampling $\sigma_{\text{psf}}$ from $[0.5,1]$ produces star images with realistic distribution patterns, as illustrated in Fig. 2. The PSF of a real star captured by a camera set to focus at infinity typically falls within this range. We also found that training the neural network with a variety of $\sigma_{\text{psf}}$ values can improve its capacity to process stars distorted by lenses or motion, which may deviate from the Gaussian PSF.
@@
where $floor()$ outputs the greatest integer less than or equal to the input value, $QE$ is the quantum efficiency of the image sensor, $FWC$ is the full well capacity of the image sensor in electrons, and $N_{bits}$ is the resolution of the ADC. A window with half $\text{half width}=floor\left(3\sigma_{\text{psf}}+0.5\right)$ centered at the centroid pixel is used to simulate the defocused star image. $V_{(u_{i},v_{i})}$ is calculated for every pixel within the window.
The pristine simulated star images are subsequently merged with noisy frames containing stray light captured by actual cameras, as illustrated in Fig. 3. Current star detection studies typically apply Gaussian noise to simulated star images [14][15]. Monakhova et al. [24] demonstrated that simplistic Gaussian or Poisson-Gaussian noise models fail to adequately represent the frequently encountered non-Gaussian, non-linear, and sensor-specific noise inherent in low-light images. Inspired by their observations, we acquire noisy dark frames from three distinct cameras by covering their lenses with lens caps. For each camera, we captured 100 frames at exposure times of 100 ms, 300 ms, 500 ms, 700 ms, and 900 ms, respectively. An additional set of 1000 frames is acquired by situating a camera within a darkroom environment and illuminating it with a flashlight from 12 distinct angles to simulate the presence of stray light. During the training image generation process, a noisy frame is randomly selected from this pool of 2500 captured frames to be fused with a noiseless star image. Random flips are applied to the noisy frame to facilitate data augmentation.
@@ 3.3 CNN Implementation Details
$L_{total}=\alpha L_{S}+\beta L_{D}$ (12)
Here, the $mean()$ function computes the average value of an array, $N$ represents the batch size, $S_{i}$ is the actual binary segmentation map, $\hat{S}_{i}$ is the estimated segmentation map with values ranging between zero and one, $D_{i}$ is the actual distance map, and $\hat{D}_{i}$ is the estimated distance map. Notation $\otimes$ denotes element-wise multiplication. For our training procedure, we set $\alpha=1$ and $\beta=2.5$ to balance the magnitude of the two loss terms.
Our dataset consists of 2,500 images for training and an additional 500 images for evaluation. CNN models undergo training for 100 epochs with a batch size set to 10 to leverage batch normalization layers effectively. We employ the Adam optimizer with a weight decay of $5\times 10^{-4}$ and an initial learning rate of $1\times 10^{-3}$. The learning rate is decreased by a factor of two every 20 epochs to facilitate better convergence. All training processes are conducted on Google Colab using a single NVIDIA V100 GPU. Our codes are available at the project website https://hongruizhao.github.io/CNNStarDetectCentroid/.
@@ 4.1 Synthetic Data Test
We employed precision, recall, and the F1 score as metrics to assess the detection performance. Precision measures the algorithm’s capability to exclude false detections, while recall quantifies its ability to identify all stars. The F1 score offers a balanced evaluation by combining both precision and recall. The mathematical expressions for these metrics are defined as
@@
In these formulas, “TP” denotes true positives, representing stars correctly identified by the star tracker. “FP” indicates false positives, and “FN” signifies false negatives, which are stars not detected by the star tracker. To evaluate centroiding accuracy, we employ the root-mean-square error (RMSE) between the actual centroid coordinates and the estimated coordinates. Additionally, we document the computational time, the total number of multiply-accumulate operations (MACs) in giga, and the number of model parameters in millions for each CNN model. All CNN models are executed on an NVIDIA RTX2060M GPU. It is worth noting that the proposed least squares centroiding process runs efficiently, requiring only 4.49 milliseconds on an AMD Ryzen 7 4800H CPU.
@@ 4.3 Static Hardware-in-the-loop Test
Figure 4: Starfield simulator allowed us to evaluate our method with real camera images.
@@
The simulator functions by displaying a set of stars on the screen from the Hipparcos Catalog, only showing stars brighter than visual magnitude 6. The camera views this visual field through the collimating lens, which is placed to focus the starfield to infinity. The images captured by the camera are then fed into the star tracking algorithms for our tests. A cover lined with light-absorbing fabric is placed over the setup to ensure minimal light leakage into the simulator. With this setup, there are some limitations. Since our screen has a limited spatial and luminance resolution, the positions and magnitude of stars are displayed with limited accuracy. The star vectors, on average, will have 12.9 arcseconds of angular distance error. The small dot pitch of our chosen monitor helps minimize this error. Based on lux measurements of the screen, we also estimate that the displayed stars have an apparent magnitude of 2.35.
With the simulator setup complete, we were able to run tests to compare our neural network and conventional baseline models. The first series involved a static starfield to measure performance, with our detection metrics being the precision, recall, and F1 scores defined in (13). To maintain consistency between tests, the simulator is calibrated by aligning the camera image center with the center of the monitor, establishing a fixed optical path. In addition, a fan is placed near the camera to provide cooling when desired, controlling the amount of thermal noise present in the camera images. Ten celestial fields of view were selected to provide test images for the algorithms. The camera integration time was fixed at 150 ms for all tests, and distortion correction was implemented using calculated OpenCV distortion coefficients. The parameters for each traditional method were kept the same for all tests; these were set to values with the best detection results for a real night sky image with 150 ms integration time. The results for the static detection are shown in Table 8.
@@
Due to the limited-resolution error and any potential distortion calibration errors, the centroiding results may not represent the true capability of the algorithms. In addition, some stars near the far edge of the screen appear stretched due to optical distortion, which makes centroiding inaccurate for those stars. To counter this, we removed outlier stars that had average angular errors of greater than 200 arcseconds. Overall, since they were all tested in the same system and with the same calibration, the consistency enables comparison between the algorithms. Specifically, the neural networks are shown to have significantly improved centroiding accuracy to the Center of Mass method, while remaining competitive with the Gaussian Grid method.
@@ 4.4 Dynamic Hardware-in-the-loop Test
Figure 5: Images depicting star pixels with zero rotation (a), $\omega$=[0.5, 0, 0] (b), and $\omega$=[1, 0, 0] (c) show that higher rotation rates produce more elongated star pixels in the starfield images.
The results for dynamic detection for $\omega$=[0.5, 0, 0], $\omega$=[0, 0, 6], and $\omega$=[1, 0, 0] are shown in Table 11. For all three rotation rates, our proposed two neural network methods perform the best with F1 scores above 98. The WITM method achieved slightly lower scores due to having more false positives. Nonetheless, it still performed significantly better than Sun et al.’s method, which struggled notably in distinguishing between true stars and false positives.
@@ 4.5 Night Sky Testing
Figure 7: Despite (a) the presence of the moon within the field-of-view or (b) severe lens flare, the proposed method leads to accurate star detection and identification.
@@ 5 Conclusion
In this article, we introduced a CNN-based method for real-time star detection and centroiding in star tracker processing. We generated training images by blending synthetic “clean” star images with authentic sensor noise and stray light. The CNN model yields a binary segmentation map for star detection and a distance map for centroiding. By incorporating distance data with pixel coordinates, centroid calculation was reformulated as a trilateration problem, which is solved using the least squares method. The chosen CNN models operates with under 300 ms latency on a low-power Google Coral TPU. We conducted synthetic image tests, hardware-in-the-loop tests, and night sky tests to thoroughly assess our method against existing approaches. Our method not only surpassed existing methods in detection and centroiding performance but also exhibited remarkable robustness across varying thermal, rotation, and stray light conditions. Moving forward, our research will progress to constructing a prototype star tracker to further validate our method in an actual flight-qualified setup. Additionally, a potential enhancement could involve training the CNN to process star trails resulting from higher rotation rates.

### v2

@@ 1 INTRODUCTION
Stray light from bright bodies or optical path dust reflections can also result in false positives. Given that the intensity of stray light often surpasses the background brightness level, conventional threshold-based star detection algorithms [5][7][8][9] are unable to discriminate false detections, undermining subsequent star identification processes. To counteract stray light, large baffles are frequently placed in front of the sensor optics [10], potentially encroaching on the limited payload space available in CubeSats. Even with a baffle, star trackers remain susceptible to blinding when directly exposed to bright bodies in their field of view [11].
@@
    We introduce an end-to-end training approach for a singular CNN, designed to concurrently handle star detection and centroiding. This represents a significant advancement over existing methods, which typically focus on either detection or centroiding in isolation.
@@
    We introduce a method for generating synthetic star images with realistic stray light and sensor noise. This approach minimizes the reliance on labor-intensive manual labeling of real star images, making our pipeline more efficient and scalable. The introduction of stray light in training data allows our method to achieve superior detection robustness.
@@
    We propose a new centroid estimation technique that leverages a distance map and trilateration, offering improved accuracy compared to conventional intensity-based methods.
  4. 4)
    We conduct a comprehensive performance evaluation of various CNN models and identifying the most suitable CNN architectures for real-time star tracker image processing.
  5. 5)
    We conduct a series of rigorous experiments with synthetic data, hardware simulator, and night sky test to evaluate the performance of our method.
@@ 3.1 Training Data Generation
where the coordinates of the real star centroid are $u_{c}$, $v_{c}$. $\sigma_{\text{psf},u}$ and $\sigma_{\text{psf},v}$ are the PSF width in width and height direction in pixels. We choose $\sigma_{\text{psf},u}=\sigma_{\text{psf},v}=\sigma_{\text{psf}}$ in our simulation. Following [6][24], we assume an ideal aberration-free lens with symmetric defoucsing pattern. In our experiments, we established that uniformly sampling $\sigma_{\text{psf}}$ from $[0.5,1]$ produces star images with realistic distribution patterns, as illustrated in Fig. 2. The PSF of a real star captured by a camera set to focus at infinity typically falls within this range. We also found that training the neural network with a variety of $\sigma_{\text{psf}}$ values can improve its capacity to process stars distorted by lens aberrations (such as coma) or motion, which may produce asymmetric star images.
@@
where $floor()$ outputs the greatest integer less than or equal to the input value, $QE$ is the quantum efficiency of the image sensor, $FWC$ is the full well capacity of the image sensor in electrons, and $N_{bits}$ is the resolution of the ADC. A window with half $\text{half width}=floor\left(3\sigma_{\text{psf}}+0.5\right)$ centered at the centroid pixel is used to simulate the defocused star image. Light contribution $V_{(u_{i},v_{i})}$ is calculated for every pixel within the window. We use this window to identify pixels belonging to a star and generate binary segmentation maps. Pixels outside this window are excluded due to their distance from the star’s center.
The pristine simulated star images are subsequently merged with noisy frames containing stray light captured by actual cameras, as illustrated in Fig. 4. Current star detection studies typically apply Gaussian noise to simulated star images [14][15]. Monakhova et al. [26] demonstrated that simplistic Gaussian or Poisson-Gaussian noise models fail to adequately represent the frequently encountered non-Gaussian, non-linear, and sensor-specific noise inherent in low-light images. Inspired by their observations, we acquire noisy dark frames from three distinct cameras by covering their lenses with lens caps. For each camera, we captured 100 frames at exposure times of 100 ms, 300 ms, 500 ms, 700 ms, and 900 ms, respectively. An additional set of 1000 frames is acquired by situating a camera within a darkroom environment and illuminating it with a flashlight from 12 distinct angles to simulate the presence of stray light. As shown in Fig. 3, our simple set up can simulate both moonlight/sunlight interference and earth-atmosphere light encountered by real satellite systems [25]. During the training image generation process, a noisy frame is randomly selected from this pool of 2500 captured frames to be fused with a noiseless star image. Random flips are applied to the noisy frame to facilitate data augmentation.
Figure 3: Our simple setup effectively simulates realistic stray light patterns when compared to real star tracker images with moonlight/sunlight (first column) and earth-atmosphere light (second column) interference[25].
@@ 3.3 CNN Implementation Details
${\color[rgb]{0,0,0}L_{total}=L_{S}+\alpha L_{D}}$ (12)
Here, the $mean()$ function computes the average value of an array, $N$ represents the batch size, $S_{i}$ is the actual binary segmentation map, $\hat{S}_{i}$ is the estimated segmentation map with values ranging between zero and one, $D_{i}$ is the actual distance map, and $\hat{D}_{i}$ is the estimated distance map. Notation $\otimes$ denotes element-wise multiplication. For our training procedure, we set $\alpha=2.5$ to balance the magnitude of the two loss terms.
Our dataset consists of 2,500 images for training and an additional 500 images for evaluation. CNN models undergo training for 100 epochs with a batch size set to 10 to leverage batch normalization layers effectively. We employ the Adam optimizer with a weight decay of $5\times 10^{-4}$ and an initial learning rate of $1\times 10^{-3}$. The learning rate is decreased by a factor of two every 20 epochs to facilitate better convergence. All training processes are conducted on Google Colab using a single NVIDIA V100 GPU.
@@ 4.1 Synthetic Data Test
We employed precision, recall, and the F1 score as metrics to assess the detection performance. Precision measures the algorithm’s capability to exclude false detections, while recall quantifies its ability to identify all stars. The F1 score offers a balanced evaluation by combining both precision and recall. Maintaining high recall is vital to ensure a sufficient number of stars are available for reliable identification, and more detected stars lead to better attitude determination accuracy in general. While modern star identification methods are robust to a reasonable number of false positives, excessive false detections can increase computational burden and potentially hinder real-time performance. Therefore, maintaining reasonable precision is still relevant.
The mathematical expressions for these metrics are defined as
@@
In these formulas, “TP” denotes true positives, representing stars correctly identified by the star tracker. “FP” indicates false positives, and “FN” signifies false negatives, which are stars not detected by the star tracker. To evaluate centroiding accuracy, we employ the root-mean-square error (RMSE) between the actual centroid coordinates and the estimated coordinates. Accurate centroid estimation is essential because it directly impacts both star identification and the calculation of precise orientation. Additionally, we document the computational time, the total number of multiply-accumulate operations (MACs) in giga, and the number of model parameters in millions for each CNN model. All CNN models are executed on an NVIDIA RTX2060M GPU. It is worth noting that the proposed least squares centroiding process runs efficiently, requiring only 4.49 milliseconds on an AMD Ryzen 7 4800H CPU.
@@ 4.2 Hardware Inference Time Test
It is important to note that this was a preliminary investigation using simple static post-training quantization technique. While it effectively demonstrated the feasibility of real-time operation on an edge AI device, the resulting detection and centroiding performance were not desirable. To fully realize the potential of our algorithm on edge AI platforms, more advanced quantization techniques, such as quantization-aware training, will be necessary. These advanced techniques are beyond the scope of this current work, but they represent an important direction for future research and development.
@@ 4.3 Static Hardware-in-the-loop Test
Figure 5: Starfield simulator allowed us to evaluate our method with real camera images. Components of the simulator are labeled accordingly.
@@
The simulator functions by displaying a set of stars on the screen from the Hipparcos Catalog limited to stars brighter than visual magnitude 6. The camera views this visual field through the collimating lens, which is placed to focus the starfield to infinity. The images captured by the camera are then fed into the star tracking algorithms for our tests. A cover lined with light-absorbing fabric is placed over the setup to ensure minimal light leakage into the simulator. With this setup, there are some limitations. Since our screen has a limited spatial and luminance resolution, the positions and magnitude of stars are displayed with limited accuracy. The star vectors, on average, will have 12.9 arcseconds of angular distance error. The small dot pitch of our chosen monitor helps minimize this error. Based on lux measurements of the screen, we also estimate that the displayed stars have an apparent magnitude of 2.35.
With the simulator setup complete, we were able to run tests to compare our neural network and conventional baseline models. The first series involved a static starfield to measure performance, with our detection metrics being the precision, recall, and F1 scores defined in (13). To maintain consistency between tests, the simulator is calibrated by aligning the camera image center with the center of the monitor, establishing a fixed optical path. In addition, a fan is placed near the camera to provide cooling to the sensor when desired, controlling the amount of thermal noise present in the camera images. The cooling from the fan reduced the sensor temperature from 29 Celsius to 24 Celsius in our tests. Ten celestial fields of view were selected to provide test images for the algorithms. The camera integration time was fixed at 150 ms for all tests, and distortion correction was implemented using calculated OpenCV distortion coefficients. The parameters for each traditional method were kept the same for all tests; these were set to values with the best detection results for a real night sky image with 150 ms integration time. The results for the static detection are shown in Table 8.
@@
Due to the limited-resolution error and any potential distortion calibration errors, the centroiding results may not represent the true capability of the algorithms. In addition, some stars near the far edge of the screen appear stretched due to optical distortion, which makes centroiding inaccurate for those stars. To counter this, we removed outlier stars that had average angular errors of greater than 200 arcseconds. Overall, since they were all tested in the same system and with the same calibration, the consistency enables comparison between the algorithms. Specifically, the neural networks are shown to have significantly improved centroiding accuracy to the Center of Mass method, while remaining competitive with the Gaussian Grid method. While the presence of close double stars has the potential to affect this accuracy, there exist only 17 pairs out of 5044 stars in the testing catalog exhibiting an angular separation of less than 0.6 arcminutes. As such, we expect minimal effect on our results due to the scarcity of these cases. Future improvements could include additional inclusions of double stars in the training data.
@@ 4.4 Dynamic Hardware-in-the-loop Test
Figure 6: Star camera images showing stars during dynamic testing (enhanced for visibility). It is seen that higher rotation rates produce more elongated star pixels in the starfield images.
The results for dynamic detection for $\omega$=[0.5, 0, 0], $\omega$=[0, 0, 6], and $\omega$=[1, 0, 0] are shown in Table 11. For all three rotation rates, our proposed two neural network methods perform the best with F1 scores above 99. The WITM method achieved slightly lower scores due to having more false negatives. Nonetheless, it still performed significantly better than Sun et al.’s method, which struggled notably in detecting faint stars.
@@ 4.5 Night Sky Testing
Figure 8: Despite (a) the presence of the moon within the field-of-view (a stray light pattern not present in our training data) or (b) severe lens flare, our method leads to accurate star detection and identification. The only fake star detected is highlighted in red. This result highlights our method’s capacity to handle diverse and unseen stray light conditions.
@@ 5 Conclusion
In this article, we introduced a CNN-based method for real-time star detection and centroiding in star tracker processing. We generated training images by blending synthetic “clean” star images with authentic sensor noise and stray light. The CNN model yields a binary segmentation map for star detection and a distance map for centroiding. By incorporating distance data with pixel coordinates, centroid calculation was reformulated as a trilateration problem, which is solved using the least squares method. We conducted synthetic image tests, hardware-in-the-loop tests, and night sky tests to thoroughly assess our method against existing approaches. Our method not only surpassed existing methods in detection and centroiding performance but also exhibited remarkable robustness across varying thermal, rotation, and stray light conditions. We found that the MobileUNet model exhibited the best performance-to-computation-time trade-off, and is the most suitable for real time star tracking applications. The model operates with under 300 ms latency on a low-power Google Coral TPU.
For future works, investigating the impact of aberrations and incorporating asymmetric patterns in the training data could further enhance performance. Another potential enhancement could involve training the CNN to process star trails resulting from higher rotation rates. To bring our algorithm closer to real edge AI applications in space, we need to explore more advanced quantization techniques to minimize performance degradation. Moving forward, our research will progress to constructing a prototype star tracker to further validate our method in an actual flight-qualified setup.